# Investigating $\Omega_{c}$ spectroscopy in two-body $\Omega_{b}$ decays

### 14 Mar 2026 v1

@@ I Introduction
A singly charmed $\Omega_{c}$ baryon is composed of three quarks, csscss, in which the two strange quarks form a correlated diquark subsystem. The orbital motion between the charm quark and the light diquark plays a central role in shaping the $\Omega_{c}$ spectrum, while possible internal orbital excitations of the diquark may also contribute. In addition, the spin of the $ss$ diquark and the spin of the charm quark provide further degrees of freedom, whose couplings with the orbital angular momentum give rise to the observed pattern of spin-parity states. Within the framework of heavy-quark effective theory, a systematic classification of the $\Omega_{c}$ spectrum has been established  Chen:2022asf ; Roberts:2007ni ; Yoshida:2015tia ; Luo:2025sns . Experimental measurements therefore play a key role in validating and refining the established $\Omega_{c}$ spectroscopy.
@@
The LHCb Collaboration has recently observed the resonant decays $\Omega_{b}^{-}\to\Omega_{c}^{**}\pi^{-},\Omega_{c}^{**}\to\Xi_{c}^{+}K^{-}$ LHCb:2021ptx , in which the states $\Omega_{c}(3000)$, $\Omega_{c}(3050)$, $\Omega_{c}(3065)$, and $\Omega_{c}(3090)$ reappear as resonance structures in the $\Xi_{c}^{+}K^{-}$ invariant mass spectrum. As a result, the existence of these states, established in the $pp$ and $e^{+}e^{-}$ collisions, is independently confirmed in weak decays. This observation opens up a new avenue for investigating the $\Omega_{c}$ spectroscopy, while simultaneously raising several important questions. First, the $\Omega_{c}(3120)$ baryon remains the only $\Omega_{c}^{**}$ state that has not been observed in this decay mode. Second, the $\Omega_{c}(3327)$ state, as well as other higher excited $\Omega_{c}$ baryons, can in principle also be produced in $\Omega_{b}^{-}$ decays; however, no corresponding signals were reported until very recently. Third, although four $\Omega_{c}^{**}$ states have been re-established in weak decays, their branching fractions have not yet been measured. In view of these issues, we consider the decays $\Omega_{b}^{-}\to\Omega_{c}^{(*)0}M$ as a suitable test ground. These channels involve the lowest-lying $1S$ baryon $\Omega_{c}^{0}$, as well as the excited $1P$- and $1D$-wave states. The final-state meson $M$, where $M$ can be $\pi$ or $\rho$, provides scalar or vector quantum numbers that couple differently to $\Omega_{c}^{(*)0}$ states.
@@ II Formalism
The $\Omega_{b}$ and the ground-state $\Omega_{c}$ baryons are both spin-1/2 fermions. In these systems, the two strange quarks form a symmetric $ss$-diquark with spin $s_{\rho}=1$, leading to the total spin $S=s_{\rho}-s_{c}=1/2$, where $s_{c}=1/2$ denotes the charm quark spin. Accordingly, their spin wave functions can be written as $\chi^{1/2}_{1/2,1}=-1/\sqrt{6}(\uparrow\downarrow\uparrow+\downarrow\uparrow\uparrow-2\uparrow\uparrow\downarrow)$ and $\chi^{1/2}_{-1/2,1}=1/\sqrt{6}(\uparrow\downarrow\downarrow+\downarrow\uparrow\downarrow-2\downarrow\downarrow\uparrow)$. As another example, $\Omega_{c}(2770)$ also contains a spin-1 $ss$ diquark ($s_{\rho}=1$) but carries total spin $S=s_{\rho}+s_{c}=3/2$. Its spin wave-function components are given by $\chi^{3/2}_{3/2,1}=\uparrow\uparrow\uparrow$, $\chi^{3/2}_{1/2,1}=(\uparrow\uparrow\downarrow+\uparrow\downarrow\uparrow+\downarrow\uparrow\uparrow)/\sqrt{3}$, $\chi^{3/2}_{-1/2,1}=1/\sqrt{3}(\uparrow\downarrow\downarrow+\downarrow\uparrow\downarrow+\downarrow\downarrow\uparrow)$, and $\chi^{3/2}_{-3/2,1}=\downarrow\downarrow\downarrow$.
@@ III Numerical Analysis
where the bracket denotes the Wigner $6j$ symbol describing the recoupling of angular momenta. In the $jj$ coupling scheme, the charmed baryons $\Omega_{c}^{(*)}$ considered in this work are labeled by $|J^{P},j_{\ell}\rangle$, as summarized in Table 3. Moreover, the subscript $\lambda$ in $|J^{P},j_{\ell}\rangle_{\lambda}$ for the $1P$-, $2S$-, and $1D$-wave states indicates that the excitation occurs in the $\lambda$ mode, while the $\rho$ mode remains in its ground state.

### 17 Mar 2026 v2

@@ I Introduction
A singly charmed $\Omega_{c}$ baryon is composed of three quarks, csscss, in which the two strange quarks form a correlated diquark subsystem. The orbital motion between the charm quark and the light diquark plays a central role in shaping the $\Omega_{c}$ spectrum, while possible internal orbital excitations of the diquark may also contribute. In addition, the spin of the ssss diquark and the spin of the charm quark provide further degrees of freedom, whose couplings with the orbital angular momentum give rise to the observed pattern of spin-parity states. Within the framework of heavy-quark effective theory, a systematic classification of the $\Omega_{c}$ spectrum has been established  Chen:2022asf ; Roberts:2007ni ; Yoshida:2015tia ; Luo:2025sns . Experimental measurements therefore play a key role in validating and refining the established $\Omega_{c}$ spectroscopy.
@@
The LHCb Collaboration has recently observed the resonant decays $\Omega_{b}^{-}\to\Omega_{c}^{**}\pi^{-},\Omega_{c}^{**}\to\Xi_{c}^{+}K^{-}$ LHCb:2021ptx , in which the states $\Omega_{c}(3000)$, $\Omega_{c}(3050)$, $\Omega_{c}(3065)$, and $\Omega_{c}(3090)$ reappear as resonance structures in the $\Xi_{c}^{+}K^{-}$ invariant mass spectrum. As a result, the existence of these states, established in the pppp and $e^{+}e^{-}$ collisions, is independently confirmed in weak decays. This observation opens up a new avenue for investigating the $\Omega_{c}$ spectroscopy, while simultaneously raising several important questions. First, the $\Omega_{c}(3120)$ baryon remains the only $\Omega_{c}^{**}$ state that has not been observed in this decay mode. Second, the $\Omega_{c}(3327)$ state, as well as other higher excited $\Omega_{c}$ baryons, can in principle also be produced in $\Omega_{b}^{-}$ decays; however, no corresponding signals were reported until very recently. Third, although four $\Omega_{c}^{**}$ states have been re-established in weak decays, their branching fractions have not yet been measured. In view of these issues, we consider the decays $\Omega_{b}^{-}\to\Omega_{c}^{(*)0}M$ as a suitable test ground. These channels involve the lowest-lying $1S$ baryon $\Omega_{c}^{0}$, as well as the excited $1P$- and $1D$-wave states. The final-state meson $M$, where $M$ can be $\pi$ or $\rho$, provides scalar or vector quantum numbers that couple differently to $\Omega_{c}^{(*)0}$ states.
@@ II Formalism
The $\Omega_{b}$ and the ground-state $\Omega_{c}$ baryons are both spin-1/2 fermions. In these systems, the two strange quarks form a symmetric ssss-diquark with spin $s_{\rho}=1$, leading to the total spin $S=s_{\rho}-s_{c}=1/2$, where $s_{c}=1/2$ denotes the charm quark spin. Accordingly, their spin wave functions can be written as $\chi^{1/2}_{1/2,1}=-1/\sqrt{6}(\uparrow\downarrow\uparrow+\downarrow\uparrow\uparrow-2\uparrow\uparrow\downarrow)$ and $\chi^{1/2}_{-1/2,1}=1/\sqrt{6}(\uparrow\downarrow\downarrow+\downarrow\uparrow\downarrow-2\downarrow\downarrow\uparrow)$. As another example, $\Omega_{c}(2770)$ also contains a spin-1 ssss diquark ($s_{\rho}=1$) but carries total spin $S=s_{\rho}+s_{c}=3/2$. Its spin wave-function components are given by $\chi^{3/2}_{3/2,1}=\uparrow\uparrow\uparrow$, $\chi^{3/2}_{1/2,1}=(\uparrow\uparrow\downarrow+\uparrow\downarrow\uparrow+\downarrow\uparrow\uparrow)/\sqrt{3}$, $\chi^{3/2}_{-1/2,1}=1/\sqrt{3}(\uparrow\downarrow\downarrow+\downarrow\uparrow\downarrow+\downarrow\downarrow\uparrow)$, and $\chi^{3/2}_{-3/2,1}=\downarrow\downarrow\downarrow$.
@@ III Numerical Analysis
where the bracket denotes the Wigner $6j$ symbol describing the recoupling of angular momenta. In the jjjj coupling scheme, the charmed baryons $\Omega_{c}^{(*)}$ considered in this work are labeled by $|J^{P},j_{\ell}\rangle$, as summarized in Table 3. Moreover, the subscript $\lambda$ in $|J^{P},j_{\ell}\rangle_{\lambda}$ for the $1P$-, $2S$-, and $1D$-wave states indicates that the excitation occurs in the $\lambda$ mode, while the $\rho$ mode remains in its ground state.